# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

### v2

@@ Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘Umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy color centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyze cantilevers with cross-sections in the range from about 0.2 to 2.6 µm2, observing a relation between cross-section and heat flux that departs from Fourier’s law predictions. We rationalize such behavior relying on first-principles simulations based on the linearized phonon Boltzmann transport equation, also discussing how fabrication-induced impurities influence conduction. Our temperature-imaging method can be applied to diamond devices of arbitrary geometry, paving the way for the exploration of unconventional, non-diffusive heat transport phenomena.
@@
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to phonon heat transport phenomena that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11, 12, 13] have been evidenced at or near room temperature, and recent theoretical work predicted that diamond should also exhibit hydrodynamic heat transport at room temperature [14]. The influence of phonon scattering at grain boundaries has been explored in polycrystalline diamond [15] and unexpected size effects were recently reported in the thermal conductivity of diamond microparticles [16], but experimental data on low-dimensional single-crystal diamond is lacking. Understanding heat transport in nanostructured diamond will be valuable in view of its applications in optomechanics [17, 18], nonlinear optics [19], quantum nanophotonics [20], as well as in heat management devices [21].
@@
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 11] and cannot provide spatially-resolved temperature measurements. Non contact optical thermometry techniques such as transient thermal gratings [22, 9], and time- or frequency-domain thermoreflectance [23, 24, 25] have been widely employed on simple planar geometries but require the deposition of a transducer material, typically a metal [26], and don’t provide a direct temperature readout [27]. Other optical techniques [28, 29], including Raman sideband thermometry [30, 31, 32, 33], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few K.
@@
In this work, we use dilute color centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. Diamond hosts optically active spin defects, the most studied being the negatively charged nitrogen vacancy (NV) center [34], which can be used as in-situ nanoscale thermometers [35, 36, 37]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [38], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [39]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [40] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [41], Si [42] and SiC [43, 44, 45], making this method widely applicable.
@@
The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm. Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source upon absorption of a 515 nm laser focused onto it (referred to as heating laser) and the bulk diamond acts as a heat sink. In the future, a UV-C laser could be used to heat diamond directly, as shown in other wide-gap materials [33]. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [46].
@@
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting (ZFS) between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [47]. The absolute value of the ZFS is found to depend on position [46], possibly due to the presence of local strain. Since our cantilevers are free-standing, we do not expect that heating induces any additional strain (except maybe very near its base). We measure the change of ZFS against a reference temperature and present in Fig. 1c the results for bulk diamond and a cantilever. Our data is fitted by the model from Ref. [48], with only two fitting parameters (dashed line in Fig. 1c).
@@
The calibration is used to convert values of ZFS shift to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [46]. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [35, 36] or on multi-frequency drive [40], the dominant sources of uncertainties in our case come from the estimates of absorbed laser power and cantilever cross-section, as detailed below.
@@
The heating laser is used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature rise as a function of heating laser power , for a readout spot located at a fixed position in the middle of the cantilever. We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [46]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the gold patch, explaining why it is slightly higher than the readout error ( K with our 3 s integration time).
@@
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, the ZFS values are measured without and with heating. To illustrate the method, the temperature profile for a cantilever of varying width is shown in Fig. 2b-c. A section of constant width shows linear temperature profile and the narrow section has a much steeper temperature gradient than the wide section, while the temperature is constant on the free-hanging side. In our experimental setting, the spatial resolution is diffraction-limited, the readout laser spot size being 0.44 µm in diameter.
@@
We can consider the case of standard diffusive heat transport to develop a picture of the expected behavior: According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity, which depends slowly on temperature. We confirmed that the Cr patch’s temperature remains close to that of the underlying diamond [46], so that convective and radiative thermal losses can be neglected. In this case, , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. For small , Fourier’s law predicts a linear temperature profile along the cantilever, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law, since it also assumes thermal conductivity to be an intrinsic, size-independent material’s property, and consequently predicts a heat flux that scales linearly with the cantilever cross-section. Instead, in diamond, has been predicted to decrease to approximately half of its bulk value at widths 0.7m [49, 50] – a much stronger reduction than other materials for this channel size[51, 52, 53, 54, 55]. We now show that a reduction of is observed when reducing the lateral size of the cantilever.
@@
We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to the difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the maximum temperature rise is comparable in all cases and remains below 40 K. The results are presented in Fig. 3, with panel (c) presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, Fourier law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross-section. Specifically, since all our cantilevers have triangular cross-section with same angle , Fourier’s law predicts . In contrast, our experiments deviate from this prediction, with a temperature gradient displaying a scaling approximately proportional to for µm.
@@
We model thermal transport in our cantilevers using the viscous heat equations (VHE), two coupled mesoscopic heat-transport equations for temperature and phonon drift velocity [14] that generalize Fourier’s laws encompassing not only heat diffusion but also fluid-like behavior [9, 10, 11, 12, 56]. The VHE can be parametrized from first principles and allow to describe how micro-structured geometries influence thermal transport with accuracy comparable to microscopic models based on the full space-dependent solution of the linearized Boltzmann transport equation (LBTE) [57, 58], at much reduced computational cost [59]. Therefore, we employ the VHE to investigate how the geometry of our cantilevers affects thermal transport, and we assess the relevance of viscous effects by comparing the VHE predictions with those obtained from Fourier’s law.
@@
We simulate the three-dimensional experimental geometry, imposing a temperature gradient K around room temperature on µm long cantilevers. Parameters entering in the VHE or Fourier’s law are computed from first-principles, accounting for third-order anharmonic phonon-phonon scattering [60, 50], phonon-impurity scattering [61], as well as for the effects of rough cantilever’s boundaries [49] (see SM [46] for details). Then, we solve the VHE or Fourier’s law and determine the heat flux from their solution. Finally, we evaluate the effective thermal conductivity as , where is the cantilever cross section, and the oriented surface element.
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [46] for details), against predictions obtained solving the VHE or Fourier’s law. In the numerical simulations, we consider two different impurity scenarios to compute transport coefficients from first principles. First, we consider the ideal case of pristine diamond cantilevers with only 0.05% 13C isotopes (dashed lines). In this idealized case, the VHE and Fourier’s law yield values having minor differences; however, both are significantly higher than the experimental values.
To address this discrepancy, we investigate theoretically the effect of additional sources of thermal resistance missing from simulations of pristine samples. In particular, FIB milling is known to cause contamination by Ga ions [62] (Fig. 4 inset) that can be as high as hundreds of ppm within a penetration depth that we estimate to be µm [46]. This penetration layer is expected to be the same for all cantilevers since we used the same FIB parameters. Consequently, Ga contamination is expected to have an increasing impact on thermal conductivity as the cantilever width decreases. We account for the influence of Ga impurities using the model of Ref. [61], considering an average concentration of Ga impurities proportional to the surface-to-volume ratio, and determine a single parameter — the concentration of Ga impurities in the penetration layer — from a fit to the experimental data [46].
Solid lines in Fig. 4 are the calculated values for within this one-parameter model, where the concentration of Ga impurities in the penetration layer obtained from the fit is ppm (a value of the same order of magnitude as those in Ref. [62]). Nitrogen and vacancy impurities are also present in the diamond with concentrations of the order of 10 ppm, which is expected to have a comparatively negligible effect on thermal conductivity [63]. A model with constant concentration of impurities over the whole cross-section fails to reproduce the experiments, especially in the steepness of the reduction of with decreasing [46]. This suggests that the main hypothesis of our model — having average concentration of Ga impurities proportional to the surface-to-volume ratio of the cantilevers — is realistic and captures the most important microscopic physics. Finally, hydrodynamic effects (quantified by the difference between green and red curves) appear practically unimportant and the deviation from Fourier’s law are mainly caused by being not an intrinsic property but a size-dependent quantity.
In conclusion, we highlight that this experimental platform consisting of NV centers in diamond nanostructures offers multiple routes for new investigations. First, the question arises of how to maximize heat conductivity in sub-micrometer diamond channels, which can be answered by achieving a finer control on the impurity levels. The effect of sidewall roughness on phonon scattering at the boundaries [64, 65] can also be probed by varying the etching method and chemical termination of the diamond surface. Second, hydrodynamic effects in heat transport could be revealed using refined geometries that enhance unambiguous signatures of hydrodynamic features, such as heat vortices [59]. Third, this research can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [66]. Overall, our results demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-Fourier regimes.

### v3

@@ Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘Umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy color centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyze cantilevers with cross-sections in the range from about 0.2 to 2.6 µm2, observing a strong reduction of the cantilevers’ conductivity as width decreases. We use first-principles simulations based on the linearized phonon Boltzmann transport equation and viscous heat equations to quantitatively predict the cantilevers’ thermal transport properties, rationalizing how the interplay between intrinsic and extrinsic phonon scattering mechanisms determines the observed non-diffusive behavior. Our temperature-imaging method paves the way for the exploration of unconventional, non-diffusive heat transport phenomena in devices and nanostructures of arbitrary geometries.
@@
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to phonon heat transport phenomena that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11, 12, 13] have been evidenced at or near room temperature. However, theoretical description of non-diffusive transport phenomena when both ballistic (extrinsic) and hydrodynamic (intrinsic) effects compete remains challenging, as it relies on phenomenological models that do not have a rigorous first-principles grounding [14, 15, 13]. In diamond, the influence of phonon scattering at grain boundaries has been explored in polycrystalline samples [16] and unexpected size effects were recently reported in the thermal conductivity of microparticles [17], but experimental data on low-dimensional single-crystal diamond is lacking. Understanding heat transport in nanostructured diamond will be valuable in view of its applications in optomechanics [18, 19], nonlinear optics [20], quantum nanophotonics [21], as well as in heat management devices [22].
@@
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 11] and cannot provide spatially-resolved temperature maps. Scanning thermal microscopy offers excellent spatial resolution, but is invasive and requires challenging calibration of the probe-sample contact thermal resistance [23, 24]. Non contact optical thermometry techniques such as transient thermal gratings [25, 9], and time- or frequency-domain thermoreflectance [26, 27, 28] have been widely employed on simple planar geometries but require the deposition of a transducer material, typically a metal [29], and don’t provide a direct temperature readout [30]. Other optical techniques [31, 32], including Raman sideband thermometry [33, 34, 35, 36], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few Kelvin.
@@
In this work, we use the spin resonance of dilute color centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. We then obtain quantitative agreement with theoretical predictions from multiscale first-principle modeling. Diamond hosts optically active spin defects, among which the negatively charged nitrogen vacancy (NV) center [37], which has been used as in-situ nanoscale thermometer [38, 39, 40], also in combination with a scanning tip [41]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [42], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [43]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [44] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [45], Si [46] and SiC [47, 48, 49], making this method widely applicable.
@@
The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm (HiQuTe Diamond). Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source upon absorption of a 515 nm laser focused onto it (referred to as heating laser) and the bulk diamond acts as a heat sink. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [50].
@@
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting (ZFS) between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [51]. We first calibrate the variation of ZFS, , against a reference bath temperature applied to the entire sample. Fig. 1c presents measured both in bulk diamond and on cantilevers. Even though the absolute value of is found to depend on position [50], possibly due to the presence of local strain, is position-independent. Indeed, since our cantilevers are free-standing, we do not expect that heating induces any additional strain (except, maybe, very near its base). Our data is fitted by the model from Ref. [52], with only two fitting parameters (dashed line in Fig. 1c).
@@
The calibration is used to convert values of to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [50]. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [38, 39] or on multi-frequency drive [44], the dominant sources of uncertainties in our case come from the estimates of absorbed laser power and cantilever cross-section, as detailed below.
@@
The heating laser is used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature rise as a function of heating laser power , for a readout spot located at a fixed position in the middle of the cantilever (see Fig. 3a,II). We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [50]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the chromium patch, explaining why it is slightly higher than the readout error ( K with our 3 s integration time).
@@
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, the ZFS values are measured without and with heating. To illustrate the method, the temperature profile for a cantilever of varying width is shown in Fig. 2c. Each section of constant width shows a linear temperature profile. The narrow section (from 4 to 10 µm) has a much steeper temperature gradient than the wide section (0 to 4 µm), while the temperature is constant on the free-hanging side ( to 0 µm). In our experimental setting, the spatial resolution is diffraction-limited, the readout laser spot size being 0.44 µm in diameter.
@@
To better understand this temperature profile, we consider the expected behavior for standard diffusive heat transport. According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity, which depends slowly on temperature. We confirmed that the Cr patch’s temperature remains close to that of the underlying diamond [50], so that convective and radiative thermal losses can be neglected. In this case, , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. For small , Fourier’s law predicts a linear temperature profile along the cantilever, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law, since it also assumes thermal conductivity to be an intrinsic, size-independent material property, and consequently predicts a heat flux that scales linearly with the cantilever cross-section. Instead, in diamond, has been predicted to decrease to approximately half of its bulk value at widths 0.7m [53, 54] – a much stronger reduction than other materials for this channel size[55, 56, 57, 58, 59]. We now show that a reduction of is observed when reducing the lateral size of the cantilever.
@@
We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to the difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the maximum temperature rise is comparable in all cases and remains below 40 K, within the calibration range of Fig. 1. The results are presented in Fig. 3, with panel (c) presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, i.e., size-independent , Fourier’s law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross-section. Specifically, since all our cantilevers have triangular cross-section with same angle , Fourier’s law predicts . In contrast, our experiments show a clear deviation from this prediction for µm. On the other hand, in the limit purely dominated by boundary scattering, thermal conductivity is expected to scale as [60], corresponding to . Our experiments are between these two regimes, indicating the necessity of accounting for both extrinsic (boundary) and intrinsic (normal and Umklapp) phonon scattering mechanisms.
@@
We model thermal transport in our cantilevers using the viscous heat equations (VHE), two coupled mesoscopic heat-transport equations for temperature and phonon drift velocity [61, 62] that generalize Fourier’s laws encompassing not only heat diffusion but also fluid-like behavior [9, 10, 11, 12, 63]. The VHE can be parametrized from first principles and allow to describe how micro-structured geometries influence thermal transport with accuracy comparable to microscopic models based on the full space-dependent solution of the linearized Boltzmann transport equation (LBTE) [64, 65], at much reduced computational cost [62]. Therefore, we employ the VHE to investigate how the geometry of our cantilevers affects thermal transport, and we assess the relevance of viscous and finite-size effects by comparing the VHE predictions with those obtained from Fourier’s law.
@@
We simulate the three-dimensional experimental geometry, imposing a temperature gradient K around room temperature on µm long cantilevers. Parameters entering in the VHE or Fourier’s law are computed from first-principles, accounting for third-order anharmonic phonon-phonon scattering [66, 54], phonon-impurity scattering [67], as well as for the effects of rough cantilever’s boundaries [53] (see SM [50] for details). Importantly, the description of finite-size effects is based on the characteristic length scale of the problem, following Casimir’s treatment of boundary scattering in small devices at low temperatures [68]111We note that this is formally analogous to Bosanquet-type regularizations for the transport coefficients of rarefied fluids [75, 62].. In this regime, where boundary scattering occurs at a rate comparable to or higher than intrinsic anharmonic phonon scattering, the appropriate length scale to use is one-half of the smallest system dimension [70]. Then, we solve the VHE or Fourier’s law and determine the heat flux from their solution. Finally, we evaluate the effective thermal conductivity as , where is the cantilever cross section, and the oriented surface element.
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [50] for details), against predictions obtained solving the VHE. We obtain good quantitative agreement between simulations and experiments. The striking deviation from diffusion, already discussed in Fig. 3, is visible as a sharp reduction of away from the bulk value in Fig. 4. A more detailed analysis of the strength of viscous hydrodynamic and ballistic effects reveals that the latter are the main contribution to the reduction of (see SM [50]). Heat transport in our cantilever is thus within the “apparent diffusion” regime [71], in the sense that it is reasonably well described by a diffusive (Fourier-like) equation with a non-intrinsic thermal conductivity (i.e., depending on extrinsic factors such as the sample size).
Fig. 4 also includes experimental data for cantilevers fabricated on distinct diamond samples, isotopically enriched in 12C and with the same NV concentration, but without relying on FIB (see SM [50]). All datasets follow the same trend, confirming in particular that possible damage induced by FIB at the cantilevers’ surface plays a negligible role in their transport properties.
In conclusion, we have demonstrated that nearly one-dimensional diamond cantilevers with sub-micrometer widths and embedded NV centers exhibit clear signatures of non-diffusive thermal transport, manifested as an effective thermal conductivity with sub-ballistic size dependence. We have quantitatively explained these observations from a first-principles multiscale approach combining the linear-response solution of the Boltzmann transport equation and the mesoscopic viscous heat equations: sub-ballistic transport arises from the interplay between intrinsic phonon interactions—including both crystal-momentum-conserving“hydrodynamic” scattering events as well as crystal-momentum-dissipating Umklapp scattering events—and extrinsic effects determined by size and geometry. This experiment sets the stage for investigating non-diffusive heat transport phenomena where intrinsic and extrinsic mechanisms coexist. First potential examples are nanostructured geometries that promote the formation of heat vortices [62], or induce thermal rectification [30]. A second example is investigating long-standing questions on the influence of sidewall roughness on phonon-boundary scattering [72, 73], since by varying the etching method and chemical termination of the diamond surface one can engineer these extrinsic parameters. Third, this experimental setup can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [74]. Overall, our results rationalize non-diffusive, sub-ballistic heat transport in diamond, and more generally demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-diffusive regimes.